A specialty store owner mixes peanuts worth 3.60 per pound with cashew nuts worth 8.40 per pound. How many pounds of each type of nut is needed to make a 12 pound mixture worth 6.80 a pound?
Answer by jorel555(1290)

Let c represent the amount of mixture of cashew nuts. Then
8.4c+3.6(12-c)=6.8(12)
8.4c+43.2-3.6c=81.6
4.8c=38.4
c=8
8 lbs of cashew nuts and 4 lbs of peanuts are needed to make the desired mixture. ☺☺☺☺